Depreciation Expense Using the Double-Declining Balance Method
The Peete Company purchased an office building for $4,500,000. The building had an estimated useful life of 25 years and an expected salvage value of $500,000.

Calculate the depreciation expense for the second year using the double-declining balance method. $
Answer #1
Depreciation Expense - 2nd Year $    331,200.00

Working:

Double Declining Balance Method = Book Value * Depreciation Rate
Depreciation Rate = Straight line depreciation rate * 200%
Straight line depreciation rate = 1/25 = 0.04
Depreciation Rate = 0.04*200% = 8%
Cost of Asset $ 4,500,000.00
Depreciation Expense - 1st Year (4500000*8%) $    360,000.00
Book Value $ 4,140,000.00
Depreciation Expense - 2nd Year (4140000*8%) $    331,200.00